OpenStudy (anonymous):

Pick a number greater than 20 and less than 150 that is not a prime number. Give all of its prime factors. How does this help you find all the possible factors of the number? (first right answer gets medal) :))

ganeshie8 (ganeshie8):

nice exercise :) prime numbers are like the building blocks of all the numbers, they can be used to generate ANY number.. pick your fav number between 20 and 150

ganeshie8 (ganeshie8):

i think so... @cheyennestaggs whats your fav number in the given range ? :)

OpenStudy (anonymous):

35 @ganeshie8

ganeshie8 (ganeshie8):

good :) what are the prime factors of 35 ?

OpenStudy (anonymous):

i think 5 and 7

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

\[\large 35 = 5^1 \times 7^1\]

ganeshie8 (ganeshie8):

since 35 is uniquely made up of these factors, you must have one of these prime numbers in all the factors of 35

ganeshie8 (ganeshie8):

you have one 5 and one 7 in the prime factorization of 35 this means you can have : 5^0 or 5^1 and 7^0 or 7^1 in a factor of 5

ganeshie8 (ganeshie8):

so total possible factors will be below : ``` 5^0 7^0 = 1 5^0 7^1 = 7 5^1 7^0 = 5 5^1 7^1 = 35 ```

ganeshie8 (ganeshie8):

So there are exactly 4 factors of 35 - let me know if something is not clear

ganeshie8 (ganeshie8):

and indeed this is true http://www.wolframalpha.com/input/?i=factors+of+35

OpenStudy (anonymous):

I got it. Thank you! :))

ganeshie8 (ganeshie8):

we can generalize this, and it may be exciting to know that there is an interesting number theoretic function which says the exact same thing : \(\large \tau \)

ganeshie8 (ganeshie8):

its called the `divisor function` wana see how to find the number of divisors a number has ? :)

OpenStudy (anonymous):

yes :)

ganeshie8 (ganeshie8):

lets make it a theorem : say \(\large N\) is a number with below prime factorization \[\large N = p_1^{e_1}p_2^{e_2}p_3^{e_3}\cdots p_r^{e_r}\] then, number of factors of \(N\) can be given as below : \[\large \text{# of factors of N} = \tau = (e_1+1)(e_2+1)(e_3+1)\cdots(e_r+1)\]

ganeshie8 (ganeshie8):

the proof is trivial - \(p_1\) can have \(0, 1, 2\cdots e_1\) powers : \(e_1 + 1\) \(p_2\) can have \(0, 1, 2\cdots e_2\) powers : \(e_2 + 1\) \(p_3\) can have \(0, 1, 2\cdots e_3\) powers : \(e_3 + 1\) \(\cdots \) \(p_r\) can have \(0, 1, 2\cdots e_r\) powers : \(e_r + 1\)

ganeshie8 (ganeshie8):

So total possible combinations will be : \((e_1+1)(e_2+1)(e_3+1)\cdots(e_r+1)\)

ganeshie8 (ganeshie8):

see if that makes more or less sense...

ganeshie8 (ganeshie8):

For example : how many factors does the number 36 have ?

ganeshie8 (ganeshie8):

well we start by writing out its prime factorization : \[\large 36 = 4\times 9 = 2^2 3^2\] so, number of factors = \(\large (2+1)(2+1) = 9\)

ganeshie8 (ganeshie8):

see if that makes more or less sense..